# MoPINNEnKF: Iterative Model Inference using generic-PINN-based ensemble Kalman filter

### v1

@@ 2.2.2 Uncertainty
In [36], the authors use the non-dominated sorting genetic algorithm (NSGA) to improve traditional stochastic gradient optimization methods (e.g., ADAM), enabling them to escape local minima more effectively. Indeed, when complete data or fully known physics information is available, PINNs combined with NSGA can effectively solve PDEs in both forward and inverse problem settings [42]. However, when data availability is limited or the data are noisy, the framework may inadvertently incorporate uncertainties originating from noisy data in forward problems. Moreover, incomplete or missing physics information can lead to wrong solutions in inverse Problems. On the other hand, NSGA-III provides an alternative method for quantifying uncertainty arising from data-driven loss functions and potentially incomplete physical models. Specifically, NSGA-III generates a set of Pareto-optimal solutions (see Section 2.2.1), which can be viewed as multiple realizations or candidate parameter configurations within the PINN framework. These solutions can be integrated into an ensemble Kalman filter (in Section 2.3), together with observational data, to approximate the posterior probability distribution. This approximation can then be iteratively utilized back to refine and correct the data loss in the PINN framework. The NSGA-III algorithm used in PINN is outlined in Algorithm 1.
1 Initialize a population $P_{0}$ consisting of $N$ individuals chosen independently and uniformly:
@@
3       Generate offspring population $Q_{t}$ of size $N$ using selection, crossover, and mutation operations.
@@
7      Apply fast non-dominated sorting [13] to partition $R_{t}$ into non-dominated fronts $\mathcal{L}_{1},\mathcal{L}_{2},\dots$ based on the multi-objective loss function defined in (3).
@@ 9      Identify the front index $i^{*}\geq 1$ such that:
13      Select subset ~$\tilde{\mathcal{L}}_{i^{*}}\subseteq\mathcal{L}_{i^{*}}$ satisfying:
@@ 2.3 Ensemble Kalman Filter (EnKF)
Since the ensemble Kalman filter (EnKF) was first developed in geophysics [16, 15], it has been widely used in atmospheric science, oceanography, and climate modeling. The EnKF involves two steps: (1). forecast (prediction) and (2). analysis (filtering). The forecast step is usually temporal propagation of state-space models:
@@
is the Kalman gain. In the ensemble Kalman filter (EnKF), both $\pi_{t\mid t-1}$ and $\pi_{t}$ are approximated by equally weighted ensembles $\{x_{t}^{(i)}\}$ and $\{\widetilde{x}_{t}^{(i)}\}$. During the update step, one first employs the prediction sample to estimate $m_{t\mid t-1}$ and $P_{t\mid t-1}$. The filter sample is then constructed by transforming the prediction sample so that its mean and covariance align with the update equations. There are several possible approaches to carry out this transformation [16, 17]. In this paper, we use the plain ensemble Kalman filter (EnKF) and the forecast step, the prior is generated using the PINNs from NSGA-III algorithm.
@@ 3 Numerical Results
In this section, we test the proposed framework, hereafter referred to as MoPINNEnKF, using two different benchmark problems: (1) the one-dimensional viscous Burgers equation and (2) the one-dimensional time-fractional mixed diffusion-wave equations (TFMDWEs). For each of these problems, we consider two scenarios: (i) the forward problem and (ii) the inverse problem. Specifically, in the inverse problem setting, we assume the diffusion coefficient is unknown for the viscous Burgers equation, while the fractional order is unknown for the TFMDWEs. Also, the observational data, $\{u^{obs}(x,t)\}$ are generated by introducing observational noise at sparse grid point, with a noise level of $\eta$. Specifically, this observational noise is drawn from a Gaussian distribution with zero mean, and its standard deviation is set equal to $\eta$ of the standard deviation of the true solution, $std(u^{truth}(x,t))$, at grid $x$. The mathematical formulation yields the following:
@@
In addition, we compare our proposed MoPINNEnKF method with two other approaches: (a) a PINN employing the traditional Adam optimizer, denoted as ADAM-PINN, and (b) a PINN employing only the NSGA-III optimizer, denoted as NSGA-III-PINN.
@@ 3.1 Burgers Equation
The one-dimensional viscous Burgers equation is a nonlinear partial differential equation frequently used as a benchmark [37, 38, 49]. The equation with Dirichlet boundary conditions, is defined on the spatial domain $\Omega=[-1,1]$ and temporal domain $[0,T]$ given by the following:
@@ 3.1.1 PINN’s Settings
In order to make a fair comparison, with different optimizers, the PINN employs a consistent deep neural network architecture comprising 8 layers with 20 neurons in each layer. During the training phase, models corresponding to each optimizer were collected until the training loss converged below a prescribed threshold $\epsilon$. Specifically, the ADAM-PINN requires $5000$ epochs to achieve convergence, whereas the NSGA-III-PINN arrives at convergence within $4$ generations. The proposed MoPINNEnKF requires of $3$ generations with $1000$ epochs to ensure the convergence of training loss.
@@ 3.1.2 Forward Problem
Table 1 shows the testing errors of ADAM-PINNs with different Gaussian noise in available observational data, with imperfect model. Because the imperfect model introduces the incorrect viscosity in Burgers equation, the PINN yields model errors. By combining noisy observation data with the model, MoPINNEnKF finds the optimal between the two. It is also noted that when the noise level is high, i.e., 80%, both the model error and observational error dominate and hence the MoPINNEnKF’s accuracy diminishes.
@@
MoPINNEnKF N/A 0.0006 0.0009 0.0014
Table 2: Mean square errors (MSEs) of three different models, ADAM-PINN, NSGA-III-PINN, MoPINNEnKF, with different noise levels in the one-dimensional Burgers Equation forward test problem.
Table 2 shows the mean square error (MSE) in the forward problem setting for (1). ADAM-PINN, (2). NSGA-III-PINN, and (3). MoPINNEnKF with different noise level of obervational data. It is shown that in all cases, MoPINNEnKF is at least one order more accurate than ADAM-PINN and NSGA-III-PINN. Specifically, when the observational is more noisy, the MoPINNEnKF stays the same level of accuracy while the other two models become less accurate. This is not surprising because both ADAM-PINN and NSGA-III-PINN use noise-contaminated observations directly in their loss functions during training, without an explicit noise-filtering mechanism. In contrast, the MoPINNEnKF framework incorporates an additional "purification" step through the Kalman filter, thereby effectively mitigating the impact of observational noise. Nonetheless, the NSGA-III-PINN exhibits slightly better robustness than ADAM-PINN, as the NSGA-III algorithm more effectively balances multiple objectives within the loss function, thereby partially mitigating the adverse effects of noisy observations.
@@
MoPINNEnKF 0.0001 $\boldsymbol{7.25\times 10^{-5}}$ 0.0020 0.0014
Table 3: Mean square errors (MSEs) of three different models, ADAM-PINN, NSGA-III-PINN, MoPINNEnKF, with $80\%$ noise levels in the one-dimensional Burgers Equation forward test problem.
Table 3 shows that the NSGA-III algorithm effectively balances the loss terms in the PINN to comparable orders of magnitude within a reasonable training time. As a result, during training, the proposed MoPINNEnKF method achieves relatively low residuals for the PDE residual loss, boundary loss, and data loss—performance that neither ADAM-PINN nor NSGA-III-PINN alone can attain.
@@
Figure 3 shows the spatiotemporal solutions and corresponding errors obtained using (1) ADAM-PINN, (2) NSGA-III-PINN, and (3) MoPINNEnKF, under different observational noise levels. Consistent with the findings summarized in Table 2, MoPINNEnKF is the most accurate model among the three approaches, with its accuracy becoming more remarkable as the noise level increases.
Figure 3: Forward problem solutions (top) and errors (bottom) for comparison of ADAM-PINN, NSGA-III-PINN and MoPINNEnKF solution for one dimensional Burgers equation; observation data noise level are chosen to be $20\%$, $50\%$, and $80\%$ respectively.
@@ 3.1.3 Inverse Problem
In the inverse problem, we assume that the diffusion term is unknown, i.e., the diffusion coefficient $\nu$ in equation (17), is not known, and its initial estimate are chosen randomly in a certain range from 00 to $1$. The diffusion coefficient is considered to be a trained parameter in the PINN framework aand we use data with varying noise levels—similar to those in the forward problem—to train each PINN model. The other neural network settings are the same as in Section 3.2.1.
Table 4 shows the estimated diffusion parameters with $L^{1}$ errors for different models. It is shown that for all different observational uncertainties, the proposed MoPINNEnKF gives the best estimate. Specifically when the observational uncertainty is small and medium, i.e., $\eta=20\%$ and $50\%$ respectively, the proposed MoPINNEnKF gives the much better estimate when comparing to PINNs with ADAM optimizer and with NSGA-III optimizer; and between the two, the later is slightly better. Yet, when the observational uncertainty is large, i.e., $\eta=80\%$, the proposed MoPINNEnKF gives similar inaccurate results as the other two. This is because when the data is noisy enough, the diffusion coefficient which is the trainable parameter together with PINN’s weights converges to the wrong that still can represent the limited noisy data. Figure 4 shows errors (left) and solutions (right) for MoPINNEnKF solution for one dimensional Burgers equation with different observation data noise level are chosen to be $20\%$, $50\%$, and $80\%$. It is consistent with the results shown in Table 4.
@@
MoPINNEnKF 0.00844 0.00526
Table 4: Comparison of ADAM-PINN, NSGA-III-PINN and the proposed MoPINNEnKF for estimating the viscosity term $\nu$ in the Burgers equation. For reference, the true value of the viscosity term is $\nu=0.01/\pi\approx 0.00318$.
Figure 4: Inverse problem errors (left) and solutions (right) for MoPINNEnKF solution for one dimensional Burgers equation; observation data noise level are chosen to be $20\%$, $50\%$, and $80\%$ respectively.
@@ 3.2 Time-fractional Mixed Diffusion-Wave Equations (TFMDWEs)
In this section, we consider the time-fractional mixed diffusion-wave equations (TFMDWEs), which generalize classical diffusion and wave equations by incorporating fractional-order time derivatives [34, 14, 35]. The TFMDWEs with Dirichlet boundary conditions are defined as follows:
@@ 3.2.1 PINN’s Settings
In order to make a fair comparison, with different optimizers, the PINN employs a consistent deep neural network architecture comprising 2 hidden layers with 50 neurons in each layer. During the training phase, models corresponding to each optimizer were collected until the training loss converged below a prescribed threshold $\epsilon$. Specifically, training ADAM-PINN requires $5000$ epochs to achieve convergence, whereas the NSGA-III-PINN reaches convergence within $4$ generations. The proposed MoPINNEnKF requires of $3$ generations with $2000$ epochs to ensure the convergence of training loss.
@@ 3.2.2 Forward Problem
MoPINNEnKF N/A 0.0850 0.0865 0.0872
Table 6: Mean absolute errors (MAEs) of three different models, ADAM-PINN, NSGA-III-PINN, MoPINNEnKF, with different noise levels in the time-fractional mixed diffusion-wave equations (TFMDWEs).
Table 6 shows the mean absolute error (MAE) in the forward problem setting for (1). ADAM-PINN, (2). NSGA-III-PINN, and (3). MoPINNEnKF with different noise level of obervational data. It is shown that in all cases, MoPINNEnKF is at least more accurate than ADAM-PINN and NSGA-III-PINN. Nonetheless, the NSGA-III-PINN exhibits slightly better robustness than ADAM-PINN, as the NSGA-III algorithm more effectively balances multiple objectives within the loss function, thereby partially mitigating the adverse effects of noisy observations. This is also reflected in the figures 68.
@@
MoPINNEnKF 0.1784 0.0166 0.0449 0.0203
Table 7: Mean square errors (MSEs) of three different models, ADAM-PINN, NSGA-III-PINN, MoPINNEnKF, with $80\%$ noise levels in the one-dimensional Burgers Equation forward test problem.
Table 7 compares the mean square error of three different models, ADAM-PINN, NSGA-III-PINN, and MoPINNEnKF, with 80% Gaussian noise in the forward problem. The results show that MoPINNEnKF consistently obtains the lowest errors across all loss components, i.e., PDE residual, boundary, and observation losses. Figure 9 shows at different time instances (top: $t=0.5$, bottom: $t=1$), the true solution (black) and forward problem solutions obtained using ADAM-PINN (red), NSGA-III-PINN (blue), and MoPINNEnKF (green) for the time-fractional mixed diffusion-wave equations (TFMDWEs) with different noise levels. The results show that MoPINNEnKF (green) outperforms both ADAM-PINN (red) and NSGA-III-PINN (blue) in terms of accuracy. Moreover, its advantage becomes even more pronounced at higher noise levels.
Figure 6: Comparison of (a) the true solution and (b) the MoPINNEnKF solution as well as the $L_{1}$ errors of (c) the ADAM-PINN and (d) the MoPINNEnKF for the time-fractional mixed diffusion-wave equations (TFMDWEs) in forward problem. Observation noise level is $20\%$.
Figure 7: Comparison of (a) the true solution and (b) the MoPINNEnKF solution as well as the $L_{1}$ errors of (c) the ADAM-PINN and (d) the MoPINNEnKF for the time-fractional mixed diffusion-wave equations (TFMDWEs) in forward problem. Observation noise level is $50\%$.
Figure 8: Comparison of (a) the true solution and (b) the MoPINNEnKF solution as well as the $L_{1}$ errors of (c) the ADAM-PINN and (d) the MoPINNEnKF for the time-fractional mixed diffusion-wave equations (TFMDWEs) in forward problem. Observation noise level is $80\%$.
@@
Figure 9: Comparison at different time instances (top: $t=0.5$, bottom: $t=0.7$) between the true solution (black) and forward problem solutions obtained using ADAM-PINN (red), NSGA-III-PINN (blue), and MoPINNEnKF (green) for the time-fractional mixed diffusion-wave equations (TFMDWEs) under varying noise levels: Panel (a) — $20\%$ noise, Panel (b) — $50\%$ noise, and Panel (c) — $80\%$ noise.

### v2

@@ 2.2.2 Uncertainty
In [lu2023nsga], the authors use the non-dominated sorting genetic algorithm (NSGA) to improve traditional stochastic gradient optimization methods (e.g., ADAM), enabling them to escape local minima more effectively. Indeed, when complete data or fully known physics information is available, PINNs combined with NSGA can effectively solve PDEs in both forward and inverse problem settings [moya2025conformalized]. However, when data availability is limited or the data are noisy, the framework may inadvertently incorporate uncertainties originating from noisy data in forward problems. Moreover, incomplete or missing physics information can lead to wrong solutions in inverse Problems. On the other hand, NSGA-III provides an alternative method for quantifying uncertainty arising from data-driven loss functions and potentially incomplete physical models. Specifically, NSGA-III generates a set of Pareto-optimal solutions (see Section 2.2.1), which can be viewed as multiple realizations or candidate parameter configurations within the PINN framework. These solutions can be integrated into an ensemble Kalman filter (in Section 2.3), together with observational data, to approximate the posterior probability distribution. This approximation can then be iteratively utilized back to refine and correct the data loss in the PINN framework. The NSGA-III algorithm used in PINN is outlined in Algorithm 1.
1Initialize a population $P_{0}$ consisting of $N$ individuals chosen independently and uniformly:
@@
3   Generate offspring population $Q_{t}$ of size NN using selection, crossover, and mutation operations.
@@
7  Apply fast non-dominated sorting [deb2002fast] to partition $R_{t}$ into non-dominated fronts $\mathcal{L}_{1},\mathcal{L}_{2},\dots$ based on the multi-objective loss function defined in (3).
@@ 9  Identify the front index $i^{*}\geq 1$ such that:
13  Select subset $\tilde{\mathcal{L}}_{i^{*}}\subseteq\mathcal{L}_{i^{*}}$ satisfying:
@@ 2.3 Ensemble Kalman Filter (EnKF)
Since the ensemble Kalman filter (EnKF) was first developed in geophysics [evensen2003ensemble, Evensen_1996], it has been widely used in atmospheric science, oceanography, and climate modeling. The EnKF involves two steps: (1). forecast (prediction) and (2). analysis (filtering). The forecast step is usually temporal propagation of state-space models:
@@
is the Kalman gain. In the ensemble Kalman filter (EnKF), both $\pi_{t\mid t-1}$ and $\pi_{t}$ are approximated by equally weighted ensembles $\{x_{t}^{(i)}\}$ and $\{\widetilde{x}_{t}^{(i)}\}$. During the update step, one first employs the prediction sample to estimate $m_{t\mid t-1}$ and $P_{t\mid t-1}$. The filter sample is then constructed by transforming the prediction sample so that its mean and covariance align with the update equations. There are several possible approaches to carry out this transformation [evensen2003ensemble, evensen2022]. In this paper, we use the plain ensemble Kalman filter (EnKF) and the forecast step, the prior is generated using the PINNs from NSGA-III algorithm.
@@ 3 Numerical Results
In this section, we test the proposed framework, hereafter referred to as iPINNER, using two different benchmark problems: (1) the one-dimensional viscous Burgers equation and (2) the one-dimensional time-fractional mixed diffusion-wave equations (TFMDWEs). For each of these problems, we consider two scenarios: (i) the forward problem and (ii) the inverse problem. Specifically, in the inverse problem setting, we assume the diffusion coefficient is unknown for the viscous Burgers equation, while the fractional order is unknown for the TFMDWEs. Also, the observational data, $\{u^{obs}(x,t)\}$ are generated by introducing observational noise at sparse grid point, with a noise level of $\eta$. Specifically, this observational noise is drawn from a Gaussian distribution with zero mean, and its standard deviation is set equal to $\eta$ of the standard deviation of the true solution, $std(u^{truth}(x,t))$, at grid $x$. The mathematical formulation yields the following:
@@
In addition, we compare our proposed iPINNER method with two other approaches: (a) a PINN employing the traditional Adam optimizer, denoted as ADAM-PINN, and (b) a PINN employing only the NSGA-III optimizer, denoted as NSGA-III-PINN.
@@ 3.1 Burgers Equation
The one-dimensional viscous Burgers equation is a nonlinear partial differential equation frequently used as a benchmark [lu2021deepxde, lu2021physics, raissi2019physics]. The equation with Dirichlet boundary conditions is defined on the spatial domain $\Omega=[-1,1]$ and temporal domain $[0,T]$ given by the following:
@@ 3.1.1 PINN’s Settings
In order to make a fair comparison, with different optimizers, the PINN employs a consistent deep neural network architecture comprising 8 layers with 20 neurons in each layer. During the training phase, models corresponding to each optimizer were collected until the training loss converged below a prescribed threshold $\epsilon$. Specifically, the ADAM-PINN requires $5000$ epochs to achieve convergence, whereas the NSGA-III-PINN arrives at convergence within $4$ generations. The proposed iPINNER requires of $3$ generations with $1000$ epochs to ensure the convergence of training loss.
Training and Testing Data
The training data for the Burgers equation consist of three main components: Initial Condition (IC) points, Boundary Condition (BC) points, and Collocation Points. The IC points are sampled in space at the initial time $t=0$, while the BC points are sampled in time along the spatial boundaries (x=1x=-1 and $x=1$). The collocation points are randomly selected from the interior of the spatiotemporal domain, with 100 points used in this study. The testing data are sampled over the domain $\Omega\times[0,T]$ with $\Omega=[0,1]$. The spatial mesh size is of $\Delta x=1/100$ and a temporal step size is $\Delta t=0.01$.
Error criteria
We evaluate models using the mean squared error (MSE) between the benchmark solutions and the predictions of different models.
Remark 1.
The iPINNER employs an ensemble strategy, which can be computationally expensive. This overhead, however, can be alleviated through parallelization, as each ensemble member is trained independently. Furthermore, the incorporation of the evolutionary multi-objective optimization algorithm (NSGA-III) accelerates the convergence of each ensemble member relative to a standard PINN. As a result, while the computational cost of iPINNER is higher, it is not prohibitive. Table 1 in revised manuscript (also attached below) list the wall time for training PINN and the proposed iPINNER. The wall time is recorded until the model reaches a convergence plateau. In addition, we performed a sensitivity analysis to compare the computational cost with respect to the ensemble size. The results, summarized in Table 2, show that the optimal ensemble size is approximately $N=8$, which provides a balance between accuracy and computational efficiency. All experiments were performed on a linux machine with an NVIDIA H100 GPU (80G).
Table 1: Wall time of training PINN and iPINNER.
Method Wall Time (s) MSE
Standard PINN 205 0.0036
iPINNER 385 0.0014
Table 2: Sensitivity analysis for the ensemble size (NN). The table shows the trade-off between the final Mean Squared Error (MSE) and the total wall-clock time. The chosen value, $N=8$, offers the best balance.
Ensemble Size ($N$) Final MSE Wall-Clock Time (s)
4 0.0019 298
6 0.0016 336
8 0.0014 385
10 0.0014 415
@@ 3.1.2 Forward Problem
Table 3 shows the testing errors of ADAM-PINNs with different Gaussian noise in available observational data, with imperfect model. Because the imperfect model introduces the incorrect viscosity in Burgers equation, the PINN yields model errors. By combining noisy observation data with the model, iPINNER finds the optimal between the two. It is also noted that when the noise level is high, i.e., 80%, both the model error and observational error dominate and hence the iPINNER’s accuracy diminishes.
@@
iPINNER N/A 0.0006 0.0009 0.0014
Table 4: Mean square errors (MSEs) of three different models, ADAM-PINN, NSGA-III-PINN, iPINNER, with different noise levels in the one-dimensional Burgers Equation forward test problem.
Table 4 shows the mean square error (MSE) in the forward problem setting for (1). ADAM-PINN, (2). NSGA-III-PINN, and (3). iPINNER with different noise level of obervational data. It is shown that, when the observational noise level is $20\%$ and $50\%$, iPINNER are at least one order more accurate than ADAM-PINN and NSGA-III-PINN. Even when the observational data become substantially noisier ($80\%$ noise level), it remains at least twice as accurate as the other two methods. This is not surprising because both ADAM-PINN and NSGA-III-PINN use noise-contaminated observations directly in their loss functions during training, without an explicit noise-filtering mechanism. In contrast, the iPINNER framework incorporates an additional "purification" step through the Kalman filter, thereby effectively mitigating the impact of observational noise. Nonetheless, the NSGA-III-PINN exhibits slightly better robustness than ADAM-PINN, as the NSGA-III algorithm more effectively balances multiple objectives within the loss function, thereby partially mitigating the adverse effects of noisy observations.
@@
iPINNER 0.0001 $\boldsymbol{7.25\times 10^{-5}}$ 0.0020 0.0014
Table 5: Mean square errors (MSEs) of three different models, ADAM-PINN, NSGA-III-PINN, iPINNER, with $80\%$ noise levels in the one-dimensional Burgers Equation forward test problem.
Table 5 shows that the NSGA-III algorithm effectively balances the loss terms in the PINN to comparable orders of magnitude within a reasonable training time. As a result, during training, the proposed iPINNER method achieves relatively low residuals for the PDE residual loss, boundary loss, and data loss—performance that neither ADAM-PINN nor NSGA-III-PINN alone can attain.
@@
Figure 3 shows the spatiotemporal solutions and corresponding errors obtained using (1) ADAM-PINN, (2) NSGA-III-PINN, and (3) iPINNER, under different observational noise levels. Consistent with the findings summarized in Table 4, iPINNER is the most accurate model among the three approaches, with its accuracy becoming more remarkable as the noise level increases.
Figure 3: Forward problem solutions (top) and errors (bottom) for comparison of ADAM-PINN, NSGA-III-PINN and iPINNER solution for one dimensional Burgers equation; observation data noise level are chosen to be $20\%$, $50\%$, and $80\%$ respectively.
@@ 3.1.3 Inverse Problem
In the inverse problem, we assume that the diffusion term is unknown, i.e., the diffusion coefficient $\nu$ in equation (17), is not known, and its initial estimate are chosen randomly in a certain range from 0 to $1$. The diffusion coefficient is considered to be a trained parameter in the PINN framework and we use data with varying noise levels—similar to those in the forward problem—to train each PINN model. The other neural network settings are the same as in Section 3.2.1.
Table 6 shows the estimated diffusion parameters with $L^{1}$ errors for different models. It is shown that for all different observational uncertainties, the proposed iPINNER gives the best estimate. Specifically when the observational uncertainty is small and medium, i.e., $\eta=20\%$ and $50\%$ respectively, the proposed iPINNER gives the much better estimate when comparing to PINNs with ADAM optimizer and with NSGA-III optimizer; and between the two, the later is slightly better. Yet, when the observational uncertainty is large, i.e., $\eta=80\%$, the proposed iPINNER gives similar inaccurate results as the other two. This is because when the data is noisy enough, the diffusion coefficient which is the trainable parameter together with PINN’s weights converges to the wrong that still can represent the limited noisy data. Figure 4 shows errors (left) and solutions (right) for iPINNER solution for one dimensional Burgers equation with different observation data noise level are chosen to be $20\%$, $50\%$, and $80\%$. It is consistent with the results shown in Table 6.
@@
iPINNER 0.00844 0.00526
Table 6: Comparison of ADAM-PINN, NSGA-III-PINN and the proposed iPINNER for estimating the viscosity term $\nu$ in the Burgers equation. For reference, the true value of the viscosity term is $\nu=0.01/\pi\approx 0.00318$.
Figure 4: Inverse problem errors (left) and solutions (right) for iPINNER solution for one dimensional Burgers equation; observation data noise level are chosen to be $20\%$, $50\%$, and $80\%$ respectively.
@@ 3.2 Time-fractional Mixed Diffusion-Wave Equations (TFMDWEs)
In this section, we consider the time-fractional mixed diffusion-wave equations (TFMDWEs), which generalize classical diffusion and wave equations by incorporating fractional-order time derivatives [liu2019alternating, du2021temporal, lu2025fpinn]. The TFMDWEs with Dirichlet boundary conditions are defined as follows:
@@ 3.2.1 PINN’s Settings
In order to make a fair comparison, with different optimizers, the PINN employs a consistent deep neural network architecture comprising 2 hidden layers with 50 neurons in each layer. During the training phase, models corresponding to each optimizer were collected until the training loss converged below a prescribed threshold $\epsilon$. Specifically, training ADAM-PINN requires $5000$ epochs to achieve convergence, whereas the NSGA-III-PINN reaches convergence within $4$ generations. The proposed iPINNER requires of $3$ generations with $2000$ epochs to ensure the convergence of training loss.
@@ 3.2.2 Forward Problem
iPINNER N/A 0.0850 0.0865 0.0872
Table 8: Mean absolute errors (MAEs) of three different models, ADAM-PINN, NSGA-III-PINN, iPINNER, with different noise levels in the time-fractional mixed diffusion-wave equations (TFMDWEs).
Table 8 shows the mean absolute error (MAE) in the forward problem setting for (1). ADAM-PINN, (2). NSGA-III-PINN, and (3). iPINNER with different noise level of observational data. It is shown that in all cases, iPINNER is at least more accurate than ADAM-PINN and NSGA-III-PINN. Nonetheless, the NSGA-III-PINN exhibits slightly better robustness than ADAM-PINN, as the NSGA-III algorithm more effectively balances multiple objectives within the loss function, thereby partially mitigating the adverse effects of noisy observations. This is also reflected in the figures 68.
@@
iPINNER 0.1784 0.0166 0.0449 0.0203
Table 9: Mean square errors (MSEs) of three different models, ADAM-PINN, NSGA-III-PINN, iPINNER, with $80\%$ noise levels in the one-dimensional Burgers Equation forward test problem.
Table 9 compares the mean square error of three different models, ADAM-PINN, NSGA-III-PINN, and iPINNER, with 80% Gaussian noise in the forward problem. The results show that iPINNER consistently obtains the lowest errors across all loss components, i.e., PDE residual, boundary, and observation losses. Figure 9 shows at different time instances (top: $t=0.5$, bottom: $t=1$), the true solution (black) and forward problem solutions obtained using ADAM-PINN (red), NSGA-III-PINN (blue), and iPINNER (green) for the time-fractional mixed diffusion-wave equations (TFMDWEs) with different noise levels. The results show that iPINNER (green) outperforms both ADAM-PINN (red) and NSGA-III-PINN (blue) in terms of accuracy. Moreover, its advantage becomes even more pronounced at higher noise levels.
Figure 6: Comparison of (a) the true solution and (b) the iPINNER solution as well as the $L_{1}$ errors of (c) the ADAM-PINN and (d) the iPINNER for the time-fractional mixed diffusion-wave equations (TFMDWEs) in forward problem. Observation noise level is $20\%$.
Figure 7: Comparison of (a) the true solution and (b) the iPINNER solution as well as the $L_{1}$ errors of (c) the ADAM-PINN and (d) the iPINNER for the time-fractional mixed diffusion-wave equations (TFMDWEs) in forward problem. Observation noise level is $50\%$.
Figure 8: Comparison of (a) the true solution and (b) the iPINNER solution as well as the $L_{1}$ errors of (c) the ADAM-PINN and (d) the iPINNER for the time-fractional mixed diffusion-wave equations (TFMDWEs) in forward problem. Observation noise level is $80\%$.
@@
Figure 9: Comparison at different time instances (top: $t=0.5$, bottom: $t=0.7$) between the true solution (black) and forward problem solutions obtained using ADAM-PINN (red), NSGA-III-PINN (blue), and iPINNER (green) for the time-fractional mixed diffusion-wave equations (TFMDWEs) under varying noise levels: Panel (a) — $20\%$ noise, Panel (b) — $50\%$ noise, and Panel (c) — $80\%$ noise.
In Figure 10, we show the finite element method (FEM) results with different mesh sizes, which demonstrate good agreement for the forward solution compared with our iPINNER. However, FEM is primarily restricted to solving forward problems. While FEM itself does not require data, with available observational data, it requires additional data assimilation techniques. On the other hand iPINNER can simultaneously addresses both forward and inverse problem settings with available noisy data.
Figure 10: Comparison of finite element method (FEM) solution with different mesh size for the time-fractional mixed diffusion-wave equations (TFMDWEs) in forward problem.